Rewrite each of the following expressions in factored form.
a. 4x+32=
b. 7x^2-18x=
c. 3x^2+15x=

Answer by Theo(13342)

a. 4x + 32 = 4 * (x + 8)

b. 7x^2 - 18x = x * (7x - 18)

c. 3x^2 + 15x = 3x * (x + 5)